### v2

# 3D Visual Illusion Depth Estimation

Abstract
3D visual illusion is a perceptual phenomenon where a two-dimensional plane is manipulated to simulate three-dimensional spatial relationships, making a flat artwork or object look three-dimensional in the human visual system. In this paper, we reveal that the machine visual system is also seriously fooled by 3D visual illusion, including monocular and binocular depth estimation. In order to explore and analyze the impact of 3D visual illusion on depth estimation, we collect a large dataset containing almost 3k scenes and 200k images to train and evaluate SOTA monocular and binocular depth estimation methods. We also propose a robust depth estimation framework that uses common sense from a vision-language model to adaptively select reliable depth from binocular disparity and monocular depth. Experiments show that SOTA monocular, binocular, and multi-view depth estimation approaches are all fooled by various 3D visual illusions, while our method achieves SOTA performance. https://github.com/YaoChengTang/3D-Visual-Illusion-Depth-Estimation
1 Introduction
Depth estimation is a task of recovering 3D geometry of a given image or image sequence. It is a long-standing and unresolved vision problem with a lot of researches, like monocular depth estimation, stereo matching, and multi-view reconstruction. These works have achieved impressive perfromance in common scenes, almost can rival that of human perception. Despite the common scenes, there are also many challenging scenes that make a flat artwork or object look three-dimensional in the visual system, like 3D video advertisements and transparent glass, as shown in Figure 1. These challenging scenes contain 3D visual illusion, which will result in wrong depth perception and seriously affect the downstream applications, casuing safety-critical risks in AR/VR and robotics.
In this paper, we present a 3D-Visual-Illusion dataset to explore the impact of 3D visual illusion on depth estimation. The dataset contains 5 kinds of illusions: inpainting illusion (e.g., inpainting on a wall/floor), picture illusion (e.g., picture printed/drawn on a paper), replay illusion (e.g., video replayed on different screens), holography illusion, and mirror illusion (e.g., virtual image on reflective objects). The dataset is collected in almost 3k scenarios with 200k images, across from indoors to outdoors, from small objects to large street views. We build the dataset using virtual data and real-world data. The virtual data is generated from web-source videos and generative models using two kind of data generation pipelines. The real-world data is collected using a stereo camera and a solid-state LiDAR depth camera. The evalution of SOTA depth estimation models on our dataset show that the monocular priors learned from large-scale data is easily fooled by illusion textures, like pictures and replayed videos, while the matching results fail at special materials, like transparent glass and mirrors.
Figure 1: The visualization of 3D visual illusion.
Inspired by the strong generalization ability of vision-language models (VLMs) across diverse tasks, we introduce a VLM-driven monocular-stereo fusion model to address the challenges posed by 3D visual illusions. The proposed model leverages the commonsense knowledge from VLM to assess the reliability of monocular and binocualr depth in different regions, enabling effective depth fusion. The model consists of a dual-branch prediction network and a VLM-based fusion network. The dual-branch prediction network takes a rectified image pair as input and uses two branches to predict monocular depth and binocular disparity simultaneously. The VLM-based fusion network first use a pre-trained large vision model to extract features from monocular depth, binocular disparity, and RGB image. The features are further transformed into a shared embedding space with language prompt using a large language model. The embedding features are then used to generate confidence map via flow matching. The confidence map is used to guide the alignment of mnocular depth from affine-invariant space to metric space. The aligned monocular depth is then fused with the binocular disparity to generate the final depth map. The experiments on our dataset and Booster dataset show that the proposed method achieves SOTA performance on different 3D visual illusions.
2 Related Work
2.1 Stere Matching
Stereo matching is a pixel-wise labeling task that relies on dense correspondence between a pair of images. The SOTA methods are either GRU-based iterative methods or Transformer-based methods. The former methods predict the disparity update and iteratively approximate the GT value in a GRU framework [26, 23, 16, 45, 5, 10]. They have achieved great performance in both benchmark and zero-shot generalization testing. Teh latter methods use Transformer to learn matching and predict the disaprity map [24, 12, 40, 41, 38, 22, 36]. The Transformer-based methods achieve superior performance by learning from a large-scale data. In this paper, we collect a comprehensive, large-scale dataset to thoroughly investigate and evaluate the impact of 3D visual illusions on matching methods. Our findings reveal that stereo matching methods are highly susceptible to various illusions, while our method leverage common sense from a large vision-language model to help rectify these illusions.
2.2 Monocular Depth Estimation
Monocular depth estimation is a pixel-wise regression task based on single image. Recent deep learning methods leverage diffusion models [17, 8, 47] or Transformers [31, 2, 44, 42] to extract depth-related features, in both supervised and self-supervised settings [9, 11, 19]. Despite their impressive generalization performance across diverse scenes, these methods fundamentally rely on monocular cues, which are susceptible to 3D visual illusions, much like the human visual system. In this work, we introduce a large-scale benchmark to evaluate state-of-the-art monocular depth models under such illusions. Our results show that existing methods are consistently misled by these challenging patterns. To address this, we propose leveraging matching-based depth cues as a complementary information to enhance monocular depth estimation.
2.3 Large Vision-Language Model
The large vision-language model (VLM) injects common sense from billions of textual data to support vision understanding and generation [46]. It presents great power in various task, like visual question answering, image generation, and navigation. The methods of these tasks mainly adapt a pre-trained VLM to specific datasets to preserve the generalization ability, while promoting the understanding of specific tasks. To further facilitate the training of VLM in downstream tasks, a lot of methods explore the different finetuning strategy, like Prompt [4], Adapter [14], LoRA [15], and LST [33]. In this paper, we propose to use VLM to predict the confidence of the disparity map to recover the metric version of monocular depth. The common sense from large VLM is beneficial for the confidence estimation in various complex scenes.
3 3D Visual Illusion Dataset
We construct a 3D-Visual-Illusion dataset to explore the challenging of 3D visual illusion in the depth estimation tasks. The 3D-Visual-Illusion dataset is constructed in almost 3k scenarios with 178912 frames used for training and 617 frames used for testing. The dataset contains images with various resolutions with the maximum resolution of $1080\times 1920$ across from indoors to outdoors, from small objects to large street inpainting. It contains 5 kinds of illusions: inpainting illusion (e.g., inpainting on a wall/floor), picture illusion (e.g., picture printed/drawn on a paper), replay illusion (e.g., video replayed on different screen), holography illusion, and mirror illusion (e.g., virtual image on reflective objects). The 3D-Visual-Illusion dataset is generated/captured in the virtual world and the real world respectively. We will introduce the details of construction for the virtual data subset and the real data subset in the following contents.
3.1 Virtual Data
We collect a large amount of video data from websites and text-to-video generative models. We take the videos as left image sequence and generate disaprity maps and right images.
Video Collection  We adopt two distinct data collection strategies for the first four types of illusions and for mirror illusions. For inpainting, picture, replay, and holography illusions, we crawl 5,226 web videos (over 52M frames) using keyword-based search. We then use a large vision-language model Qwen2-VL-72B [1, 37] to automaticly delete bad frames, reducing the dataset to 4,519 videos (1.4M frames). Then, we manually remove blurry or occluded frames, resulting in 1,384 high-quality videos (236K frames). For mirror illusions, due to the rarity of mirror-related keywords, mirror illusion videos are hard to find from the web. Instead, we employ generative models (Sora [28], Kling [21], and HunyuanVideo [20]) to generate videos. Prompts are initially generated with ChatGPT and iteratively refined. After manual filtering, we collect 234 high-quality mirror illusion videos containing 2,382 frames.
Depth Generation  After collecting videos from both web and generative models, we generate depth for each frame using the pipelines shown in Figure 2 and 3. We use different data generation pipelines for the two kinds of data due to the following reasons: (1) Web-sourced data mainly use fixed cameras, offering limited viewpoints and making accurate scene reconstruction challenging. (2) Videos from generative models are used for mirror illusions, requiring the geometry of the mirror world for right image generation. However, monocular depth estimation may predict the geometry of either the mirror surface or the reflected scene, making it difficult to control.
For web videos, we use the pre-trained DepthAnything V2 [43] to predict inverse depth, interpreted as disparity under unknown camera parameters. However, the generated disparity in 3D illusion regions is completely wrong. To address this, we introduce a neighboring support region as a reference to correct the disparity of illusion regions by assuming they are located on the same plane. Segmentation masks for both illusion and support regions are obtained using SAM2 [32]. Due to the limitations of automatic mode in detecting illusions, we manually annotate all frames using SAM2’s click mode, filtering out redundant or imperceptible cases. This yields 176,530 high-quality annotated frames.
Figure 2: The data generation pipeline for web-sourced data.
Figure 3: The data generation pipeline for videos from generative models.
After obtaining the mask of illusion and support regions, a plane can be fitted using the points within a support region. In the standard 3D camera coordinate system $(X,Y,Z)$, the general form of a plane with parameters $(\alpha,\beta,\gamma,\delta)$ is $\alpha\cdot x+\beta\cdot y+\gamma\cdot z+\delta=0$. Given the relationship between image coordinate $(u,v)$, disparity $d$, and $(X,Y,Z)$:
$(u,v,d)=\frac{1}{z}(x,y,B)\cdot(f_{x},f_{y},\frac{f_{x}+f_{y}}{2})+(c_{x},c_{y},0),$
the planar structures in 3D space $(X,Y,Z)$ remain planar in the disparity space $(u,v,d)$ with plane parameters $(\alpha,\beta,\delta,\gamma)$: $\alpha\cdot u+\beta\cdot v+\delta\cdot d+\gamma=0$. This is crucial for our generation, as it allows us to fit a 3D plane using the disparity values under unknown camera baseline and intrinsics, without converting disparity into depth. Given a set of $N$ points from the support region, $\{(u_{i},v_{i},d_{i})\}_{i=1}^{N}$, the goal of plane fitting becomes a least squares fitting problem about parameters $(\alpha,\beta,\delta,\gamma)$. To mitigate the impact of noise during fitting, we adopt RANSAC [7] for robust plane fitting (For more details, please refer to the supplemental materials). The fitted plane is then used to rectify disparity values within illusion regions. After obtaining the rectified disparity map, we further denoise the disparity to accquire smooth boundaries between support regions and their surroundings.
For videos from generative model, we reconstruct the entire scene using InstantSplat [6], which first derives geometry from DUSt3R [38] and generates novel views using GS [18]. We extract disparity information from the DUSt3R geometry and further refine it through a series of post-processing steps, similar to aforementioned semantic segmentation, RANSAC plane fitting, and disaprity denoising.
Right Image Generation  The right-view images for generative-model videos are directly rendered using Gaussian Splatting (GS). For web-sourced videos, right views are synthesized by warping the left images using monocular disparity. Due to scale ambiguity, we estimate an optimal scale factor $s$ via binary search, terminating when most warped pixels fall within the valid image width:
$\tilde{s}=\arg\min_{s}\left|\frac{\sum_{x}\mathbbm{1}(0\leq u-s\cdot d_{(u,v)}<W)}{N}-\tau\right|,$ (1)
where $\mathbbm{1}(\cdot)$ is the indicator function, $d$ is the disparity, $(u,v)$ are pixel coordinates, $W$ is the image width, $N$ is the number of pixels, and $\tau$ is the target ratio of valid pixels. We use the scaled disparity $\tilde{d}=\tilde{s}\cdot d$ to warp the left image accordingly. In cases where multiple source pixels are warped to the same target location due to occlusions, we retain the one with the largest disparity to maintain consistency:
$I_{r}(u^{\prime},v)$ $=I_{l}(u^{*},v),$ (2)
$u^{*}=\arg\max_{u}\{d_{(u,v)}$ $\mid u-\tilde{d}_{(u,v)}=u^{\prime}\}.$
To address holes after warping, we apply an image inpainting method [34] to produce visually complete right-view images. The full generation algorithm is detailed in the supplemental materials.
3.2 Real-world Data
In addition to virtual data, we collect real-world data comprising 72 scenes and 617 frames. The setup includes a stereo camera (ZED Mini) and a LiDAR-based depth camera (Realsense L515). Details of the setup are provided in the supplemental materials. To ensure accurate alignment, the two cameras are rigidly mounted, and their relative pose is calibrated using a checkerboard. Based on the calibration, the L515 depth map is warped to the ZED left camera to construct GT depth.
Due to the lower resolution of the L515, direct pixel-wise warping to the higher-resolution ZED image will result in sparse depth maps and incorrect projections, particularly in occluded regions where background depths may overwrite foreground pixels. To address this, we first densify the L515 point cloud by upsampling its depth map $\mathcal{Z}_{L}$ via nearest-neighbor interpolation and proportionally scaling its intrinsic matrix $K_{L}$.
After densifying the point cloud, image coordinates from the L515, $(U_{L},V_{L})$, are first projected into the 3D camera coordinate $(X_{L},Y_{L},Z_{L})$ using the L515 depth map. These 3D points are then transformed to the ZED left camera’s coordinate system $(X_{Z},Y_{Z},Z_{Z})$. Finally, they are projected onto the ZED left image coordinates $(U_{Z},V_{Z})$:
$[X_{L},Y_{L},Z_{L}]$ $=Z_{L}\cdot K_{L}^{-1}\cdot[U_{L},V_{L},1],$ (3)
$[X_{Z},Y_{Z},Z_{Z}]$ $=R\cdot[X_{L},Y_{L},Z_{L}]+T,$
$[U_{Z},V_{Z},1]$ $=K_{Z}\cdot[X_{Z}/Z_{Z},Y_{Z}/Z_{Z},1].$
Here, $R\in\mathbb{R}^{3\times 3}$ and $T\in\mathbb{R}^{3\times 1}$ denote the rotation and translation matrices between the L515 and the ZED cameras, both obtained via calibration. $K_{L}\in\mathbb{R}^{3\times 3}$ and $K_{Z}\in\mathbb{R}^{3\times 3}$ represent the intrinsic matrices of the L515 and the ZED left camera, respectively.
In the projected coordinates $(U_{Z},V_{Z})$, multiple 3D points $P_{m}$ may map to the same pixel due to slanted surfaces or occlusions. To resolve this, we apply Z-buffering to retain the point with the minimum depth for the ZED depth map $\mathcal{Z}_{Z}$:
$\mathcal{Z}_{Z}(u_{Z},v_{Z})$ $=z_{Z}^{*},$ (4)
$z_{Z}^{*}=\min_{(u_{Z}^{{}^{\prime}},v_{Z}^{{}^{\prime}},z_{Z}^{{}^{\prime}})\in P_{m}}\{Z_{Z}^{{}^{\prime}}\mid$ $(u_{Z}^{{}^{\prime}},v_{Z}^{{}^{\prime}})=(u_{Z},v_{Z})\},$
Although upsampling greatly densifying point cloud, projecting from a lower-resolution to a higher-resolution space may still introduce small holes. To address this, we apply connected component analysis to identify missing regions and fill them using image inpainting [35]. The inpainted depth values are further smoothed to ensure seamless transitions with surrounding areas.
Figure 4: The pipeline of VLM-driven binocular and monocular disparity fusion model.
Furthermore, to identify unreliable depth values, the depth map projected onto the ZED image is reprojected back to the L515 coordinate:
$[X_{Z\rightarrow L},Y_{Z\rightarrow L},Z_{Z\rightarrow L}]$ $=R^{-1}\cdot(Z_{Z}\cdot K_{Z}^{-1}\cdot[U_{Z},V_{Z},1]-T),$ (5)
$[U_{Z\rightarrow L},V_{Z\rightarrow L},1]$ $=\frac{1}{Z_{Z\rightarrow L}}\cdot K_{L}[X_{Z\rightarrow L},Y_{Z\rightarrow L},Z_{Z\rightarrow L}].$
The reprojected pixels that correspond to invalid depth values or exhibit large depth differences with their ZED counterparts are marked as unreliable:
$\mathcal{Z}_{Z}(u_{Z},v_{Z})$ =$=\begin{cases}0&\text{if }\mathcal{Z}_{L}(u_{Z\rightarrow L},v_{Z\rightarrow L})==0\ or\\ &\quad\left|z_{Z\rightarrow L}-\mathcal{Z}_{L}(u_{Z\rightarrow L},v_{Z\rightarrow L})\right|>\tau,\\ \mathcal{Z}_{z}(u_{z},v_{z})&\text{otherwise}.\end{cases}$ (6)
Here, $\tau$ is a manually defined threshold. To further refine depth quality, median filtering is applied to suppress noise and remove outlier points from the point cloud. Finally, the depth map $\mathcal{Z}_{Z}$ is converted into disparity map $D$:
$D=B\cdot F/\mathcal{Z}_{Z},$ (7)
where $B$ denotes the baseline between the ZED’s stereo cameras, and $F$ is the focal length of the ZED camera. The entire algorithm is detailed in the supplemental materials.
4 VLM-Driven Monocular-Stereo Fusion Model
We draw inspiration from the strong generalization ability of vision-language models (VLMs) across diverse tasks to address the 3D-visual-illusion challenge. We propose a VLM-driven monocular-stereo fusion model that leverages commonsense knowledge to assess the reliability of monocular and stereo depth in different regions, enabling effective depth fusion. As illustrated in Figure 4, our model contains two components: a dual-branch prediction network and a VLM-based fusion network.
4.1 Dual-Branch Prediction Network
As shown in Figure 4, our dual-branch prediction network comprises a binocular disparity estimation branch and a monocular depth estimation branch. The stereo branch adopts an iterative optimization framework, extracting features from rectified image pairs and constructing a cost volume via dot product. Starting from an initial disparity of zero, a GRU-based module iteratively refines the disparity map. The monocular branch takes the left image as input and uses the fixed-weight DepthAnything V2 [43] to predict affine-invariant inverse depth, representing normalized relative disparity. We also extract features from the penultimate layer of DepthAnything V2 and adapt them via learnable convolutional layers. These adapted features serve as left-view context to guide disparity refinement in the stereo branch.
4.2 VLM-Based Fusion Network
As shown in Figure 4, the VLM-based fusion network contains three main stages: the VLM prediction stage, the confidence map generation stage, and the global fusion stage.
VLM Prediction Stage  In this stage, we utilize the pretrained QwenVL2-7B model [37, 1]. The visual prompts consist of the left-view RGB image, binocular disparity map, and monocular disparity map. Due to the complexity and variability of monocular texture-based ambiguities, which are hard to describe precisely in text, the textual prompts are designed primarily from a binocular perspective. Specifically, they focus on reflective, transparent, or other material types that typically cause stereo matching failures. By leveraging the vision-language model’s general reasoning capabilities, we extract embedding features that help distinguish between monocular and binocular depth cues under language guidance.
Confidence Map Generation Stage  This stage aims to transform the embedding features back into image space to generate a confidence map. Inspired by the flow-matching framework Flux [25, 29], we learn a guided path flow from Gaussian noise to a complex confidence distribution:
$y_{k_{i+1}}=y_{k_{i}}+\Delta K\cdot v_{k_{i}}(y_{k_{i}},c_{e}),$ (8)
where $k$ is uniformly sampled in the interval $[0,1]$, $\Delta K=1/K$, and $K$ is the total number of steps. $y_{0}$ is sampled from a prior Gaussian distribution, $y_{k_{i}}$ is the intermediate state at $i$-th sampling, $c_{e}$ is the conditional embedding, and $v_{t_{k}}$ denotes the predicted velocity field at time $t_{k}$ under conditions $(y_{t_{k}},c_{e})$.
Specifically, the textual prompts are first mapped into the embedding space. The image and text embeddings are concatenated to form $c_{e}$, which is combined with the intermediate state $y_{t_{k}}$ and passed through a stack of Transformers. Cross-attention is used to inject conditional information, and the velocity field $v_{t_{k}}$ is predicted. After multiple iterations via Equation 8, the final state $y_{1}$ is tensorized into a 2D format and decoded into image space using a variational autoencoder. The resulting feature is then concatenated with the cost volume and passed through convolutional layers to predict the confidence map $I_{c}$:
$I_{c}=\sigma(\mathcal{F}_{c}([\mathcal{G}_{c}(y^{\prime}_{1}),V(D_{s})])),$ (9)
where $\sigma$ is the sigmoid function, $\mathcal{F}_{c}$ denotes convolution, $\mathcal{G}_{c}$ is the VAE decoder, $y^{\prime}_{1}$ is the tensorized form of $y_{1}$, and $V(D_{s})$ is the cost volume sampled around binocualr disparity $D_{s}$.
Global Fusion Stage  Gloabal fusion first aligns monocualr disparity $D_{M}$ to the absolute disparity space and then fuse the aligned monocualr disparity $\tilde{D}_{m}$ with binocualr disaprity $D_{s}$. The alignment is achieved by affine transformation parameters $s_{m},t_{m}$:
$\tilde{D}_{m}$ $=s_{m}\cdot D_{m}+t_{m},$ (10)
$s_{m},t_{m}=\arg\min_{s_{m},t_{m}}$ $\sum_{(u,v)}(s_{m}\cdot D_{m}(u,v)+t_{m}-D_{s}(u,v))^{2}.$
$s_{m}$ and $t_{m}$ are learned through convolutions on the concatenation of the monocular disparity $D_{m}$ and the binocular disparity $D_{s}$. Since $s_{m}$ and $t_{m}$ on low-confident regions is unrealible, we further refine the parameters by pooling $s_{m}$ and $t_{m}$ of the high-confident neighbors. After acquiring refined parameters $s_{m}$ and $t_{m}$, we compute the aligned monocular disparity $\tilde{D}_{m}$ using Equation 10. $\tilde{D}_{m}$, $D_{s}$, and $I_{c}$ are then concatenated and passed through convolutions and upsampling layers to generate the final high-resolution disparity map.
5 Experiment
We first train the models on the SceneFlow dataset [27] and then fine-tune them on our 3D-Visual-Illusion dataset. The finetuned models are then tested on the 3D-Visual-Illusion test set and the Booster training set [30]. We compare with monocular depth estimation methods (DepthAnything V2 [43], Marigold [17], Metric3D [44], and DepthPro [3]), multi-view large model methods (Dust3R [38] and VGGT [36]), and stereo matching methods (RAFT-Stereo [26], Selective-IGEV [39], and MochaStereo [5]). For more training details, loss, evaluation metrics, and prompts, please refer to the supplementary material.
Table 1: Evaluation comparison in disparity and depth space on the 3D-Visual-Illusion dataset. ‘+ align’ indicates alignment using globally shared affine parameters computed from ground truth.
Method Disparity Space Depth Space
EPE $\downarrow$ bad2 $\downarrow$ bad3 $\downarrow$ bad5 $\downarrow$ AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$uparrow
DepthAnything V2 [43] 5.81 61.45 43.18 30.57 0.14 0.15 92.86
Metric3D [44] 12.46 94.11 91.14 82.05 0.34 0.29 48.97
DepthAnything V2 metric [43] 16.24 92.53 87.43 75.25 0.52 0.39 48.75
DepthPro [3] 12.26 87.08 80.60 62.43 0.28 0.25 65.92
Marigold [17] 21.16 65.67 59.67 53.19 0.45 0.37 63.65
DepthAnything V2 metric [43] + align 5.23 56.82 45.50 28.89 0.17 0.15 93.70
Metric3D [44] + align 5.70 66.26 50.92 40.43 0.17 0.17 94.80
DepthPro [3] + align 4.36 44.98 34.98 24.70 0.09 0.10 93.83
Dust3R [38] 6.74 52.89 45.31 36.61 0.25 0.22 87.09
VGGT [36] 6.16 53.32 44.89 37.20 0.13 0.12 78.46
RAFT-Stereo [26] 1.62 24.32 13.20 2.97 0.04 0.06 99.18
Selective-RAFT [39] 1.58 23.46 12.65 2.57 0.03 0.07 99.60
Selective-IGEV [39] 1.67 24.06 13.11 2.99 0.04 0.10 99.26
MochaStereo [5] 1.75 25.49 14.11 3.54 0.04 0.11 98.76
StereoAnything [13] 2.41 29.00 16.15 6.54 0.11 0.32 96.23
ours 1.77 26.72 15.73 3.60 0.03 0.08 99.60
Table 2: Comparison on Booster dataset in quarter resolution and balanced set. Metrics are reported separately for all regions (All), transparent regions (Trans), and nontransparent regions (NonTrans).
Method All Trans NonTrans
EPE bad2 bad3 bad5 EPE bad2 bad3 bad5 EPE bad2 bad3 bad5
DepthAnything V2 [43] 3.16 48.83 36.98 21.22 7.19 77.98 69.02 50.91 2.91 47.25 35.09 19.25
Metric3D [44] 35.55 99.70 99.32 97.73 41.55 99.37 98.93 97.91 34.89 99.71 99.32 97.59
DepthAnything V2 metric [43] 21.55 94.28 91.37 84.21 28.42 93.04 90.72 86.78 20.94 94.25 91.22 83.84
DepthPro [3] 24.44 92.98 90.23 84.25 25.65 92.98 88.90 83.05 24.14 92.91 90.16 84.19
Marigold [17] 5.99 57.90 47.13 32.63 8.46 76.33 65.90 51.52 5.72 56.79 45.87 31.26
DepthAnything V2 metric + align 5.71 62.70 48.94 32.18 12.72 77.24 68.46 54.70 5.45 62.05 48.17 31.36
Metric3D + align 3.09 43.05 29.65 16.85 8.72 76.87 64.68 47.62 2.76 41.28 27.91 15.22
DepthPro + align 4.02 53.76 40.30 23.81 6.02 64.25 55.19 40.39 3.96 53.25 39.61 23.12
Dust3R [38] 3.70 48.57 34.16 19.53 8.69 73.40 64.03 50.18 3.34 47.21 32.56 17.93
VGGT [36] 3.70 34.05 23.44 14.58 10.78 72.22 65.12 55.83 3.32 32.27 21.34 12.24
RAFT-Stereo [26] 4.08 17.61 14.87 12.17 9.55 67.84 59.43 47.46 3.13 13.10 10.70 8.63
Selective-RAFT [39] 4.05 19.48 16.64 13.57 10.08 70.02 61.79 49.64 2.92 14.94 12.38 9.93
Selective-IGEV [39] 4.52 19.23 16.51 13.84 9.22 67.00 58.99 47.21 3.52 14.69 12.28 10.20
MochaStereo [5] 3.79 16.77 14.24 11.77 9.18 66.64 58.10 45.78 2.82 12.25 10.11 8.30
StereoAnything [13] 4.36 24.13 19.20 14.50 10.54 73.48 63.53 49.37 3.29 20.09 15.37 11.08
ours 2.43 13.84 9.98 6.91 7.32 56.77 47.83 36.45 1.76 10.06 6.54 4.08
5.1 3D-Visual-Illusion Dataset
Tables 1 report a comprehensive comparison on the 3D-Visual-Illusion dataset. In this dataset, we find monocular depth estimation methods struggle significantly in inpainting, replay, and picture illusions, resulting in bad error rate in both disparity and depth spaces. Even using GT to align the monocular relative depth with absolute depth, the performance is till far bad than ours. Compared to recent foundation model (Dust3R [38] and VGGT [36]), our method shows clear advantages. Although these models combine monocular priors with explicit matching constraints, they exhibit a strong monocular bias on illusion-rich scenes, posing risks for downstream tasks like robotics and AR. Compared to stereo matching methods relying on explicit matching, our method achieves comparable results, showing that our method keep the matching constrains. We also provide visualization in Figures 5. Our method shows superior robustness to 3D visual illusions.
5.2 Booster Datset
Booster dataset [30] contains many objects with specular and transparent surfaces. We use this dataset to evalu1ate the genralization performance in mirror illusions. As shown in Table 2, our method achieves the best performance across all regions, including transparent and non-transparent areas. This demonstrates the effectiveness of our method in handling complex visual illusions, particularly in challenging scenarios involving reflections and refractions. We also provide visualization in Figure 6. Our method shows superior robustness to mirror illusions.
Figure 5: Visualization on 3D-Visual-Illusion dataset.
Figure 6: Visualization on Booster dataset
Table 3: Abalation study on the Booster dataset. MF: Monocular Feature, PF: Post Fusion, APF: Adaptive Post Fusion, SF: Stereo Fusion, VLM: Vision-Language Model.
MF PF APF SF VLM epe bad2 bad3 bad4 bad5 bad6 bad7
15.11 80.38 72.35 66.06 61.32 57.04 52.97
8.36 69.89 61.01 53.50 47.47 42.16 37.43
9.25 68.46 59.03 51.48 45.86 40.29 35.60
9.59 72.77 61.95 52.90 46.31 40.28 35.12
10.40 81.94 67.57 57.82 50.17 44.81 39.69
7.32 56.77 47.83 41.48 36.45 32.28 28.75
5.3 Ablation Study
We perform ablation studies on the Booster dataset to evaluate the contribution of each module. As shown in Table 3, the baseline stereo-only model performs poorly, confirming that binocular cues alone are insufficient in challenging material regions. Adding monocular depth via simple post fusion (PF) significantly improves generalization. Incorporating high-level monocular features (MF + PF) offers further gains, highlighting the benefits of feature-level fusion. Dual GRU fusion (APF) brings marginal improvement and, in some cases, introduces noise due to redundant or conflicting updates. Using static monocular priors repeatedly in stereo GRU (SF) worsens performance, showing that naive reuse of uncertain priors can be harmful. Finally, our full VLM-based fusion yields the best results, improving accuracy by over 10 points in bad2, demonstrating its strong reasoning ability and robustness in visually ambiguous regions.
6 Conclusion
In this paper, we introduce the 3D-Visual-Illusion dataset, a large-scale benchmark designed to evaluate depth estimation models under the influence of 3D visual illusions. The dataset covers a diverse range of illusion types and scenes, spanning both indoor and outdoor environments, and includes both virtual and real-world data. We also propose a baseline method: a VLM-driven monocular-binocular fusion model that leverages commonsense reasoning from a vision-language model to assess the reliability of monocular and binocular depth cues across different regions. Experiment results show that our dataset reveals key limitations in existing monocular and large-scale matching methods. Furthermore, our fusion framework significantly improves depth estimation performance in illusion-affected regions.
Limitation and Discussion  Our virtual data generation pipeline relies on manual semantic segmentation, which is labor-intensive and time-consuming. In addition, our real-world data currently covers only a subset of illusion types, specifically inpainting, picture, and replay illusions. We plan to expand the dataset to include a broader range of illusions and more diverse scenes in future work. Our model train the VLM via LoRA fine-tuning. While effective, this setup does not fully exploit the potential of foundation models. Future work will explore the learning of foundation model on illusion data to better leverage VLMs for depth estimation under visual illusions.